OpenStudy (anonymous):

Complex coeff Cn? x(t,a) = 0, -1/2

OpenStudy (anonymous):

To calculate Cn The formula is \[Cn = 1/T \int\limits_{<T>}^{} x(t)e ^{-i \omega nt}dt\]

OpenStudy (anonymous):

i dont even understand x(t,a) = 0, -1/2<t< -a/2 1/a, -a/2<t<a/2 0, a/2<t<1/2

OpenStudy (anonymous):

i'll help upto my knowledge...but pls verify with other too... now x(t,a) means that x is a function that depends on t and a, that is, if the value of either t or a change, the value of x will change as well... next x(t,a)=0 when t is in between the range -1/2 and -a/2..and also in the range a/2 and 1/2 x(t,a)= 1/a when t is in the range -a/2 to +a/2..this is can be understood by diagram...